### v3

# SmileyLlama: Modifying Large Language Models for Directed Chemical Space Exploration

Abstract
Here we show that a general-purpose large language model (LLM) chatbot, Llama-3.1-8B-Instruct, can be transformed via supervised fine-tuning of engineered prompts into a chemical language model (CLM), SmileyLlama, for molecule generation. We benchmark SmileyLlama by comparing it to CLMs trained from scratch on large amounts of ChEMBL data for their ability to generate valid and novel drug-like molecules. We also use direct preference optimization to both improve SmileyLlama’s adherence to a prompt and to generate molecules within the iMiner reinforcement learning framework to predict new drug molecules with optimized 3D conformations and high binding affinity to drug targets, illustrated with the SARS-Cov-2 Main Protease. This overall framework allows a LLM to speak directly as a CLM which can generate molecules with user-specified properties, rather than acting only as a chatbot with knowledge of chemistry or as a helpful virtual assistant. While our dataset and analyses are geared toward drug discovery, this general procedure can be extended to other chemical applications such as chemical synthesis.
1Kenneth S. Pitzer Theory Center and Department of Chemistry,
University of California, Berkeley, CA, 94720 USA
2Department of Computer Science, University of Wisconsin–Madison, Madison, WI, 53706
3Department of Molecular Medicine, The Herbert Wertheim UF Scripps Institute for Biomedical Innovation and Technology, 130 Scripps Way, Jupiter, FL, 33458 USA
4Departments of Bioengineering and Chemical and Biomolecular Engineering,
University of California, Berkeley, CA, 94720 USA
5Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA
$\dagger$ thg@berkeley.edu
1 Introduction
Chemical Language Models (CLMs)1 trained on string representations such as Simplified Molecular-Input Line-Entry System (SMILES) 2 and SELF-referencIng Embedded Strings (SELFIES) 3 have emerged as a useful tool for de novo generation of molecules, best exemplified by molecules relevant to pharmaceutical applications and drug discovery4, 5, 6, 7, 8, 9. Nearly all CLMs for molecular generation have been trained from scratch on large quantities of data such as ChEMBL10 and ZINC11 and using different model architectures, including variational autoencoders12, recurrent neural networks (RNNs) long short term memory (LSTM) cells 13, 14, generative pre-trained transformers (GPTs)15, 16 and structured state space sequence (S4) models17, 18. In addition, CLMs have achieved advances in chemical generation tasks through further downstream optimization of molecules with additional training or different model frameworks.7, 8, 19, 20, 21 A recent such example is the iMiner algorithm that operates on SELFIES and uses deep reinforcement learning (RL) with real-time 3D docking to generate molecules with high potential for inhibiting protein targets.21
Language models are statistical models of probability distributions of units of language, and can be adapted to generate meaningful text by sampling from these distributions.22 The most recent advancement of language models have resulted from the training of scaled-up transformers 23 on massive amounts of data, resulting in the creation of Large Language Models (LLMs) such as GPT4 and Llama.24, 25, 26. While LLMs were originally created for applications in natural language processing, in recent years there have also been efforts to adapt LLMs to chemistry. Some studies have accessed frontier LLMs for the purpose of assisting research in the form of virtual lab members, or even performing research autonomously.27, 28, 29 Other studies have focused on training LLM architectures on chemical data which can translate between natural and chemical languages30, 31 such as finding a molecule that is a close match to a ground truth molecule from which the description is assessed by fingerprint tanimoto similarity.32 Beyond chemical dictionary lookups, LLMs have also been used to perform mutation and crossover for an evolutionary algorithm to explore chemical space.33 There have also been studies of LLMs to modify SMILES strings to change the properties of the molecules that they represent.34, 35, 36, 37 Others have taken inspiration from LLMs to design CLMs with new capabilities, such as the ability to respond to prompts through a transformer-based architecture37, 38. However, to our knowledge, no CLM derived from a pre-trained general purpose LLM foundation model has reached the performance exhibited by modern CLMs that are trained from scratch with chemical data.
Here, we demonstrate that an open-weight LLM, Meta-Llama-3.1-8B-Instruct (which we will refer to as “Llama” for the remainder of the paper) 26, can be converted into a CLM for generative tasks in drug discovery. Using supervised fine-tuning (SFT) and Direct Preference Optimization (DPO)39 of the pre-trained Llama with SMILES strings derived from ChEMBL, SmileyLlama generates drug-like molecules with desirable properties specified in a user-defined prompt with relevance to medicinal chemistry. SmileyLlama not only uses the ability of LLMs to have their behavior guided by prompts, but also their ability to serve as a foundation model which can be fine-tuned to specialize in various text-generation tasks for molecule generation that we show can match or exceed the performance of modern CLMs trained directly on chemical data. We further demonstrate that SmileyLlama greatly improves the reinforcement learning component of iMiner to more efficiently explore chemical space to create molecules optimized for 3D binding to target proteins, illustrated with the SARS-Cov-2 (SARS2) main protease (MPro).40 While our dataset and subsequent analyses are created with drug discovery in mind as a downstream application, this general procedure can be extended to other chemical applications such as chemical synthesis41.
2 Methods
2.1 Supervised Fine-Tuning and Direct Preference Optimization
In order to steer the outputs of the pre-trained Llama model 26 for drug molecule generation we first use supervised fine-tuning (SFT), in which the weights of Llama are further optimized on generic drug-like properties (Section 3.1). We constructed our fine-tuning dataset with the SMILES strings of approximately 2 million molecules from ChEMBL Dataset v33 10. For each molecule in our dataset, we randomly picked a number of properties of pharmaceutical interests to calculate using RDKit 42. These properties include ranges of hydrogen-bond (H-bond) donors and acceptors, molecular weight (MW), log of the partition coefficient (logP), number of rotatable bonds, the fraction of $sp^{3}$ carbons (i.e., the number of $sp^{3}$ carbons divided by the total number of carbon atoms in the molecule, termed $Fsp^{3}$), the presence/absence of macrocycles, the presence/absence of covalent warhead-related SMARTS patterns, the presence/absence of at least one undesirable SMARTS pattern, a SMILES string representation of a BRICS single-pass substructure and the specification of a chemical formula.
For numerical properties, we choose ranges that are relevant for medicinal chemistry. In many medicinal applications, molecules must have suitable characteristics of drugs related not just to their chemical architecture but also to relevant biological phenomena (e.g., oral absorption, metabolism, distribution). For example, molecules invented in the early stages of drug design are typically smaller in size and lipophilicity than their follow-on analogs and optimized drugs, so a user can specify that a model obeys the Rule of Three—a common rule of thumb for early small molecule drug design—when generating molecules43. As another example, we choose the TPSA ranges based on those which tend to be orally bioavailable or able to pass through the placenta or blood-brain-barrier44. If a drug need not meet these criteria (e.g., an injectable drug for a peripheral target may not require as specific a TPSA range) then a user interfacing with SmielyLlama should be able to adjust the range criterion or eliminate it. Further information about the specifics of these properties and the ranges we choose to specify during training can be found in the Supplementary Information in Section 1.
After calculating and picking these properties for each SMILES string, we construct a prompt containing values of these properties, with the “correct” completion being the SMILES string that these properties were calculated from. For example, we used a prompt with a system instruction of You love and excel at generating SMILES strings of drug-like molecules and a user instruction of the form Output a SMILES string for a drug like molecule with the following properties: if properties are specified, or Output a SMILES string for a drug like molecule: if no properties are specified. This trains our model to generate molecules that have properties specified (or not) in the prompt. We structure the prompts used for SFT so that, during inference, the properties of generated molecules can be specified in the prompt. This lets users avoid having to downselect the vast majority of generated molecules for having the correct characteristics—instead, users can simply prompt SmileyLlama to provide molecules with the characteristics they desire. See Figure 1 and Algorithm 1 in the Supporting Information for further elaboration of SFT training.
Figure 1: A visualization of the SFT workflow for Smiley-Llama. Given the Llama-3.1-8B-Instruct model 26, we used prompt-response pairs consisting of calculated molecular properties and completed SMILES strings to fine-tune Llama on SMILES strings completions, yielding SmileyLlama. Crucially, we construct the prompt for each example using properties calculated from the correct response (a SMILES string from ChEMBLv33).
We also use Direct Preference Optimization (DPO)39 which also updates the weights of Llama such that the language model is more likely to produce strings designated as "winners" and less likely to produce strings designated as "losers". We use DPO to further reinforce our model’s ability to robustly generate molecules with properties specified in the prompt for more specific task-oriented goals (Sections 3.2 and 3.3). Algorithmically we prompt our SFT model to generate molecules with a given property, and sample several SMILES strings and use RDKit42 to assess whether they have properties in line with what the prompt requested. We then pair up molecules which correctly followed the prompt with those that don’t as winners and losers, respectively, then use a single epoch of DPO to improve the model’s results. See Algorithm 2 in SI for pseudocode of this scoring and pairing procedure.
Finally we consider DPO for generating unique and valid ligands which bind to a specific protein target, Sars2-Mpro using the iMiner framework, the details of which have been published elsewhere 21. During our optimization process, we prompt SmileyLlama to provide molecules with a “High SARS2PRO” score. After every 2000 distinct molecules are generated per iteration, we score each molecule and randomly pair it with another. The molecule with the higher score is designated as the "winner" and the other as the "loser", and we construct a dataset with eight such pairings for each molecule, and then optimize SmileyLlama using DPO with respect to this dataset. We use the iMiner scoring function of three times the docking score plus a custom drug-likeness function to score all molecules per iteration. Invalid molecules were assigned a score of -10. This process is described in Algorithm S3 in the Supplementary Information.
2.2 Additional Training Details
We performed both SFT and DPO on Llama using the Axolotl Package 26, 45. For both SFT and DPO, we use the Low-Rank Adaptation (LoRA) applied to the linear layers of the model and FlashAttention with an Adam optimizer, cross-entropy loss, and a cosine learning rate scheduler with a maximum learning rate of $2\times 10^{-5}$.46, 47, 48, 49 All prompts were formatted according to the Alpaca instruction format 50, 51. Additional parameters for our training are a LoRA rank of 32, a LoRA alpha of 16, a LoRA dropout of 5%, and 10 warmup steps. For SFT, we trained for 1 epoch using a batch size of 64 on a single 4xA40 node for approximately 1 day with a validation set $\sim$5the size of the original. We also note that we randomized the SMILES string representation of each molecule, and we tokenized all SMILES strings with the Llama3 tokenizer26 when interfacing with SmileyLlama.
3 Results
3.1 Comparison of SmileyLlama from SFT with other CLMs
To test the generative ability of SmileyLlama compared with other existing CLMs across a range of architectures, we used the GuacaMol suite 52 to benchmark the validity, uniqueness, and novelty of the molecules as shown in Table 1. Additionally, KL divergence and Frechet Chemnet Distance 53 based on the GuacaMol definition (FCDGuac) are used to analyze the distributional shifts from the ChEMBL training data for drug-like molecules more specifically.
Table 1: GuacaMol benchmarks comparing Llama, SmileyLlama, and some common CLM architectures trained on ChEMBL. The model benchmarks include valid chemical molecules, uniqueness and novelty with respect to the training set, and distribution similarity evaluated using KL divergence and Frechet ChemNet distance based on the GaucaMol definition ${}_{Guac}=exp(-0.2*FCD)$. All benchmarks were performed with 10,000 samples at a temeperature $T=$1.0 and a maximum of 256 new tokens for Llama and SmileyLlama. Due to the proprietary nature of Llama’s training data and the SMILES contained therein, we cannot assess novelty, KL-Divergence or FCDGuac for Llama’s zero-shot performance.
Benchmark Validity Uniqueness Novelty KL div FCDGuac
VGAE-MCTS54 1.000 1.000 1.000 0.659 0.009
LSTM18 0.983 0.999 0.848 0.993 0.901
GPT18 0.915 1.000 0.978 0.977 0.826
S4 18 0.971 0.997 0.961 0.994 0.853
Llama zero-shot 0.693 0.458 N/A N/A N/A
Llama twenty-shot 0.463 1.000 0.952 0.921 0.081
SmileyLlama 0.960 1.000 0.988 0.970 0.553
We first analyze the ability of Llama to produce molecules, relying only on its pre-trained knowledge (zero-shot), or by providing it with one or more examples from the ChEMBL database in the formulated prompt (Table 1 and Supplementary Table S1). We find that even without SFT or examples provided in a prompt the LLM is able to produce some valid and unique SMILES strings, but is generally poor even with variations in hyperparameters such as temperature ($T$) compared to other state-of-the-art CLMs. In Table 1 it is seen that SFT significantly improves SmileyLlama’s ability to generate drug-like molecules as good as the CLMs, although FCDGuac is fairly low compared to CLMs.
That said, we still find that SmileyLlama generates very good agreement with ChEMBL quantities across a diverse property set as seen in Figure 2. The UMAP visualization in Figure 2a finds that while there are small relative differences in densities, SmileyLlama generates molecules in every well-represented region of the chemical space of ChEMBL. Likewise, the distribution of molecular properties of interest to medicinal chemistry are in very good agreement between the SmileyLlama-generated molecules and ChEMBL molecules as shown in Figure 2b. Hence there are no obvious chemical regions of ChEMBL that are inaccessible to SmileyLlama. Furthermore, it has recently come to our attention that certain undesirable molecular scaffolds might be oversampled in CLM generative tasks.21 Supplementary Table S1 shows that although small percentages of these unviable chemical structures are present in the ChEMBL training data itself, all CLMs do not oversample these undesirable molecular scaffolds. Finally, while training was conducted at $T=1.0$, exploration of temperature used at inference on the Guacamol benchmark (Supplementary Fig. 1) suggests that this temperature is adequate for all tests described in the Results.
Figure 2: Distribution comparisons for different properties of the generated molecules from SmileyLlama (blue) with molecules from the training dataset from ChEMBL (gold). (a) UMAP visualization of a random selection of 10,000 ChEMBL molecules and 10,000 SmileyLlama-generated molecules, using 15 neighbors and a minimum distance of 0.1. (b) The molecular properties considered are fraction of $sp^{3}$ hybridized carbons and heteroatoms, number of heavy atoms, number of H-bond donors and acceptors, number of aliphatic and aromatic rings and the maximum ring size, number of rotatable bonds, quantitative estimate of drug-likelihood (QED) value55, MW, approximate log partition coefficient between octanol and water (ALOGP)56, polarizable surface area (PSA) and topological PSA 57, and the number of structural alerts 58. All benchmarks were at a temperature $T=$1.0 and a maximum of 256 new tokens.
3.2 Property Specification
In Table 2 we show the average percentage of valid, distinct SMILES strings generated for a complete panel of families of tasks. This benchmark is distinct from other conditional molecule generation benchmarks16, 32 in that we are testing SmileyLlama’s ability to robustly generate molecules with properties in value ranges rather than molecules with a property of a specific value. This is for the reason of what is of interest to medicinal chemistry; many of the "rules of thumb", such as counting the number of Lipinski violations or number of hydrogen bond donors and acceptors, are based on value ranges during chemical exploration. Additionally, LLMs tend to struggle with numbers which have many degrees of precision and must be split into several tokens59. Hence we did not represent this category in the prompt during training.
Table 2: Percentage of valid, distinct generated molecules over a more complete panel of tasks using SmileyLlama. We performed a panel of 380 tasks (see Supplementary Information for complete definition) to assess how well SmileyLlama follows directions in a prompt. For each task, we generate 1000 molecules from a prompt requesting some property, and score the result based on the proportion of molecules which are valid, distinct, and satisfy the properties requested in the prompt. Finally, we collect some tasks into families (such as those with different range values) and average their scores to produce the results shown in the table below. All benchmarks were at a temperature $T=$1.0 and a maximum of 128 new tokens to increase computational efficiency. We also note that running SFT for two epochs rather than one did not seem to greatly affect the performance on these benchmarks.
SFT DPO
Present in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0
$\leq$ k H-bond donors 96.6 % 94.4 % 99.2 % 98.1 %
$\leq$ k H-bond acceptors 96.3 % 90.7 % 98.3 % 98.4 %
$\leq$ k Molecular weight 90.1 % 84.3 % 97.8 % 98.0 %
$\leq$ k ClogP 89.2 % 85.4 % 96.8 % 97.7 %
Rotatable bonds in range 90.0 % 86.7 % 94.5 % 94.0 %
Fraction $sp^{3}$ in range 87.6 % 85.0 % 96.3 % 95.9 %
TPSA in range 95.9 % 91.1 % 98.8 % 98.4 %
No bad SMARTS 92.5 % 89.6 % 94.9 % 94.3 %
Absence of macrocycle 98.4 % 95.3 % 98.4 % 97.3 %
Absence of warhead-related SMARTS 96.4 % 94.2 % 97.3 % 95.4 %
Lipinski rule-of-five 89.0 % 81.7 % 98.7 % 97.7 %
Presence of warhead-related SMARTS 58.3 % 51.0 % 74.9 % 73.0 %
Presence of Enamine substructures 51.0 % 51.5 % 67.7 % 70.1 %
Presence of macrocycle 38.8 % 44.7 % 58.0 % 57.8 %
Absent in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0
Rule-of-three 77.5 % 62.0 % 84.8 % 94.5 %
Exactly k H-bond donors 21.4% 19.7% 30.7% 30.4%
Exactly k H-bond acceptors 19.9% 14.4% 27.0% 31.9%
Overall SmileyLlama model does very well on tasks on which it was trained through the engineered prompt. Although some of the properties were present in training, such as the Lipinski rule-of-5 and the presence of a macrocycle, both were more uncommon in the data and hence performance in these categories are less good. Similarly while there were many warhead-related SMARTS and Enamine substructures that were specified in the SFT training procedure, some were more common than others and performance is thus more middling as also seen in Table 2. Additionally, some tasks did not occur at all in the SFT training data, and yet are still parsible as valid SMILES. Finally, and as expected, SmileyLlama does poorly on tasks involving exact numerical specifications. We note that one has a choice to use SmileyLlama using lower temperatures at inference that can improve the SFT predictions as also seen in Table 2. SmileyLlama also performs well in driving not just singular molecular design tasks, but on compound tasks such as generating molecules similar to existing leads, i.e. “scaffold hopping” R-group modification, and/or structure-based design to grow molecules from ligand fragments. Figure 3a is an example of SmileyLlama model’s ability to generate molecules from all 320 substructures in the Enamine database61 that follow the Lipinski rule-of-five62, which encompasses most of the molecular properties with ranges listed in Table 2.
Figure 3: Conditional generation with SmileyLlama for fragment growth and before and after DPO compared to ChEMBL. (a) Example molecules generated by growing from one of the Enamine substructures and to satisfy Lipinski’s Rule of 5 using the prompt Output a SMILES string for a drug like molecule with the following properties: a substructure of O=C(O)c1ccc(C(F)(F)F)cc1, <= 500 MW, <=5 logP, <= 5 H-bond donors, <= 10 H-bond acceptors. (b) distribution of four properties satisfying Lipinski’s rule of five comparing ChEMBL molecules (orange) with molecules generated by SmileyLlama (blue) with the prompt Output a SMILES string for a drug like molecule with the following properties: <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 MW, <= 5 logP, compared to 1000 molecules generated by SmileyLlama with the same prompt after DPO (gray). MW and LogP distributions were estimated using a gaussian kernel density estimator (KDE).60 All results generated 1000 molecules at a temperature $T=$1.0 and a maximum of 128 new tokens.
While the SFT SmileyLlama model typically does well on tasks on which it was trained through the engineered prompt, and can still perform adequately when queried with prompts different from ones on which it was trained (Table 2), we can further optimize SmileyLlama for specific tasks using DPO. DPO’s most popular application has been in improving the responses of LLM-derived chatbots, such as the instruct-tuned Llama3 models, but it has also found use in improving the outputs of CLMs 63 and avoiding the need to separately train a reward model 39. Here the relevance of DPO provides a way to further optimize the model by pairing desirable responses with undesirable responses. The model’s weights are then updated to be more likely to produce the ‘winner’ of the pairing and less likely to produce the ‘loser’ of the pairing. We generated our dataset by simply pairing up unsuccessful attempts at generating structures with successful attempts randomly for each task in Table 2. We concatenate these pairings into a single dataset and optimize for 1 epoch using DPO. SmileyLlama optimized with DPO significantly improved adherence to the prompt across nearly all tasks as seen in Table 2 and Figure 3b. Note that while DPO does cause the model to more robustly obey the rules in the prompt, it also shifts and narrows the property distribution compared to the training set as seen in Figure 3b and appears to be largely insensitive to temperature. SmileyLlama without DPO, on the other hand, occasionally does not obey the prompt but more faithfully reproduces the distribution of properties found in a filtered ChEMBL that satisfy Lipinski’s rules. Hence, our investigation indicates that DPO cna be quite useful for some tasks, but can sometimes have unintended consequences.
3.3 3D Optimization of Drugs in Protein Active Sites with SmileyLlama
The tests performed in Sections 3.1 and 3.2 do not take advantage of the 3-dimensional structural information of a putative drug nor its shape and molecular compatibility with a target protein active site. Hence we use SmileyLlama to generate unique and valid ligands which undergo further optimization for binding to a specific protein target when embedded in the iMiner framework21. iMiner generates novel inhibitor molecules for target proteins by combining deep reinforcement learning64, 65 with real-time 3D molecular docking using AutoDock Vina66, thereby simultaneously creating chemical novelty while constraining molecules for shape and molecular compatibility with target active sites. More specifically, the iMiner generative model uses an Average Stochastic Gradient Descent Weigh-Dropped Long Short-Term Memory (AWD-LSTM)67 recurrent neural network that predicts probability of string tokens to concatenate on a molecular string representation until a complete molecule is generated. In the subsequent RL stage, 2000 molecules in each epoch (typically $\sim$ 50 epochs) are sent to AutoDock simulations in parallel, and the docking scores are circled back to the RL model to adjust its parameters so that molecules generated in the next iteration will have better scores while retaining drug-likeness.
Given the ascendancy of attention mechanisms and transformer architectures, such as those inherent in SmileyLlama, we replaced the generative AWD-LSTM component of iMiner with SmileyLlama, and replaced the iMiner optimization algorithm, Proximal Policy Optimization (PPO)68 with DPO. To validate the effectiveness of SmileyLlama in the iMiner context, we generate inhibitor molecules for the popular SARS2 MPro as an example. We chose this system because of the readily available experimental 3D structures69, 70 which provides the necessary information needed for structure-based ligand design using SmileyLlama and comparisons to other successful inhibitor molecules.
For the unconditional de novo generation case, SmileyLlama learns the user prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO which pertains to minimizing the AutoDock Vina score while maximizing the drug likeliness score ($S_{DL}$) of the original iMiner reward function21. Figure 4a compares the docking scores of the original iMiner algorithm against SmileyLlama as a function of epoch number and with number of generated molecules per iteration. We notice first an improved data efficiency compared to iMiner, in which SmileyLlama requires roughly one-quarter of the epochs to reach a similar level of improved docking score. Due to iMiner’s use of SELFIES as its molecular representation, its validity is unparalleled, however, in later epochs, diversity crashes which explains the sharpening peaks in later iterations (Figure 4) and quantified further in Supplementary Figure S2 against the Guacamol benchmarks. This simply reflects convergence in the docking score, i.e. there are fewer novel molecules as docking score reaches the highest values. By contrast, SmileyLlama maintains more diversity while greatly improving the docking score to iMiner (Figure 4a) with some degradation in validity compared to iMiner (Supplementary Figure S2) likely due to the use of the SMILES representation.
Figure 4b shows the property distributions of the final optimized set of novel molecules from SmileyLlama from the prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO. While the property distributions are satisfactory for parts of the Lipinski Rule of 5, such as number of hydrogen bond donors and
Figure 4: Comparison of the shift in docking score distributions for iMiner compared to SmileyLlama over optimization epochs as illustrated for SARS2-MPro. (a) For iMiner, in later epochs diversity crashes which explains the sharpening peaks in later iterations. SmileyLlama with DPO (SL+DPO) enforces diversity throughout the optimizations (Algorithm S3), which accounts for the broad peaks, and shows superior data efficiency relative to iMiner. (b) We compare two different user prompts: Sars2Pro and Sars2Pro+Ro5. All results were generated with 1024 SMILES at a temperature of $T=$1.0 and a maximum of 128 new tokens.
acceptors, the MW and logP results are not conforming to drug-like values. This indicates some inadequacy of the iMiner reward function, i.e. a non-optimal weighting of docking scores compared to $S_{DL}$. While a standard CLM such as iMiner would require a reweighting and/or new terms in the loss/reward function, other hyperparameter tuning, and expensive retraining, a unique advantage of SmileyLlama is that the distribution of generated molecules’ properties can be shifted using nothing more than prompt engineering, with no retraining required. Figure 4b shows that combining prompts such as Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO, <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 molecular weight, <= 5 logP , which we will refer to as the High SARS2Pro+Ro5 prompt improves properties such as MW and logP and drug-likeness scores significantly with some expected loss in high docking scores since smaller molecules make fewer intermolecular interactions.
Figure 5: SmileyLlama de novo generated molecules in the active site of SARS2 main protease. Surface rendering of the SmileyLlama generated molecules in the SARS2 Mpro canonical binding pocket. Generated by SmileyLlama after optimization with (a) the SARS2PRO prompt. (b) and (c) the SARS2Pro+Ro5 prompt. Supplementary Table S2 provides their SMILES string and docking scores, and Supplementary Figure S3 shows their docking pose, for some of the highest scoring ligands.
Figure 5 provides a set of novel molecules from SmileyLlama docked in the MPro active site with the two engineered prompts High SARS2Pro and High SARS2Pro+Ro5. Two of the higher scoring molecules resemble the variations of the perampanel drug with the trefoil structure which are tested inhibitors optimized by the Jorgensen group.71 However, unlike the molecules from their study that consistently retained the central pyridinone ring71, SmileyLlama molecules have replaced the trefoil hub with the pyrimidine functional group (see Figure 5c). Higher docking scores are found for quite different drug scaffolds (Figure 5(a,b)), but in all cases there is no significant homology match found in the Therapeutic Target Database.72 This would indicate that the generative capabilities of SmileyLlama are robust and outside of the pre-trained Llama model. Finally, the proposed drugs are synthetically accessible73, as made clear from their on average SAsimilar-$\sim$ 3 as seen in Supplementary Table S2.
3.4 SmileyLlama Outside of Chemical Language Modeling
While SFT and DPO clearly alters Llama in the creation of SmileyLlama, one might wonder whether SmileyLlama still retains the full LLM knowledge of the Llama-3.1-8B model from which it derives. We find that SmileyLlama can still converse in English if it is prompted to do so, and some sample conversations are included in Section 3 in the Supplementary Information. However, we notice that its responses are often shorter than those of Llama-3.1-8B and can contain SMILES string non-sequiturs, which seem to be more likely when dealing with prompts relating to chemistry. As a more quantitative measure of its residual abilities, we evaluate performance of SmileyLlama on the Language Model Evaluation Harness containing the MMLU GPQA, Math-Hard, and MMLU-Pro benchmarks74, 75, 76, 77, 78. Supplementary Table S3 and Figure S4 show that SmileyLlama generally performs worse on moral scenarios, and amusingly SmileyLlama also performs worse on subjects dealing with chemistry than Llama. This in part is likely due to the tendency for SmileyLlama to complete prompts relating to chemistry with a SMILES string as noted above. In addition, accuracy errors in the MMLU tests have also been noted recently79, and thus SmileyLlama’s degraded performance in chemistry may also be in part an artifact of poor testing benchmarks. Overall this result is somewhat encouraging, since it implies the possibility that LLM-derived CLMs can inherit and take advantage of the natural language processing ability of their foundation model. SmileyLlama already does this— we can steer the properties of the molecules it generates and the chemical space it explores using natural language prompts while still retaining some ability to process non-chemical natural language. However, more work will be required to view SmileyLlama as an additional capability of an LLM, as opposed to its current and primary benefit as a CLM.
Discussion and Conclusion
We have investigated the adaptation of LLMs to molecular chemistry problems that have been largely in the domain of CLMs. We find that the key advantage that LLMs have when performing as CLMs comes from prompting: information about the region of interest in chemical space can be specified in the prompt, letting the model generate chemical structures in this region. Furthermore, we showed that training an LLM on a variety of prompts allows for good performance in multiple CLM tasks, and found evidence to suggest that the LLM models can correctly generalize to scenarios with new prompts that leverage the abilities developed when trained on similar prompts. Concretely, we showed that Meta-Llama-3.1-8B-Instruct can be converted into a CLM on par with or exceeding the performance of modern CLMs7, 8, 19, 20, 21, rather than only demonstrating that LLMs can modify currently existing SMILES strings to have different properties.
Our study clarifies a few crucial points for CLMs derived from LLMs going forward. First, it is not necessary to pretrain a specialized model on chemistry-specific text to generate molecules from a text description; a much less resource-intensive SFT training run on prompt-following on a dataset of a few million molecules with a commodity LLM is enough to do that. Second, DPO provides another resource-efficient way of optimizing the model to produce molecules that score well on a targeted objective without needing in-context examples, instead relying on the generative nature of the model itself for good and bad examples. Using these optimization methods, we showed that data specific to a target protein can generate increasingly better inhibitor molecules as determined by docking scores. A surprising corollary to this is the finding that SmileyLlama can combine its knowledge gained during single-objective optimization to perform well at a task specifying multiple objectives, elicited by combining the prompts. SmileyLlama was able to generate molecules which satisfy Lipinski’s rule of 5 and a high docking score on SARS2-Mpro even though SmileyLlama was not trained on the generation of molecules with both prompts, but instead is found to generate molecules with both attributes at inference.
The prompting and optimization framework for modifying LLMs to explore chemical space broadly or to narrow the search to specific regions shown here could also be leveraged for molecular design outside of drug discovery, such as the use of SMILES for elaborating on transition metal complexes80. One could also imagine that casting a chemical problem as a linguistic construct could also allow for new applications, such as our recent work on chemical synthesis41. As with many of the fields touched by LLMs this decade, the newly opened frontier of possibility in chemistry is as vast as it is exciting.
Acknowledgments
This work was supported in part by the National Institute of Allergy and Infectious Disease grant U19-AI171954 for the drug molecule application. We thank the CPIMS program, Office of Science, Office of Basic Energy Sciences, Chemical Sciences Division of the U.S. Department of Energy under Contract DE-AC02-05CH11231 for support of the machine learning. We thank Riza Özçelik for kindly providing the retraining code for different CLMs for benchmarking and Nicole Kennedy for suggesting properties of molecules useful to medicinal chemists.
Data Availability
The data included here contains all SMILES strings we used from ChEMBL (chembl.txt), the dataset containing the BRICS substructures of these ChEMBL molecules, the dataset containing prompts and responses used for SFT to create SmileyLlama from Llama, and the dataset used to create SmileyLlama-DPO-opt from SmileyLlama. Source data for all Figures is available with this manuscript. Data used for this study can be found at https://anonymous.4open.science/r/SmileyLlama-gh-D038.
Code Availability
Code used for this study can be found at https://anonymous.4open.science/r/SmileyLlama-gh-D038/
Author Contributions Statement
J.M.C.,K.S., A.G., and T.H.G. defined goals and designed the project. J.M.C., K.S., A.G., and D.B. carried out the optimizations. J.M.C., K.S., A.G., and T.H.G. wrote the paper. All authors discussed the results and made comments and edits to the manuscript.
Competing Interests Statement
The authors declare no competing interests.
Supplementary Information
SmileyLlama: Modifying Large Language Models for Directed Chemical Space Exploration
Joseph M. Cavanagh1, Kunyang Sun1, Andrew Gritsevskiy2, Dorian Bagni1, Thomas D. Bannister3,
Teresa Head-Gordon1,4,5†
1Kenneth S. Pitzer Theory Center and Department of Chemistry,
University of California, Berkeley, CA, 94720 USA
2Department of Computer Science, University of Wisconsin–Madison, Madison, WI, 53706
3Department of Molecular Medicine, The Herbert Wertheim UF Scripps Institute for Biomedical Innovation and Technology, 130 Scripps Way, Jupiter, FL, 33458
4Departments of Bioengineering and Chemical and Biomolecular Engineering,
University of California, Berkeley, CA, 94720 USA
5Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA
$\dagger$ thg@berkeley.edu
S1 Details of properties to fine-tune
Overview of selected properties for fine-tuning. When fine-tuning Llama to generate drug-like molecules, we carefully assess various design choices and proceed with the following properties, emphasizing those that medicinal chemists would consider when proposing de novo drug molecules. We categorized and summarized all 12 properties into 4 subgroups as follows.
  • Physiochemical properties. Absorption, distribution, metabolism, and excretion (ADME) are the crucial criteria to quantify the localization and concentration of drug molecules within the body after administration. As a result, we build on the list of properties proposed in the classical Lipinski’s rule of 5 62 with some modern additionstpsa, WEI20201839 to generate drug-like molecules that could demonstrate decent ADME.
    • Number of Hydrogen Bond Donors (#HBD)
    • Number of Hydrogen Bond Acceptors (#HBA)
    • Molecular Weight (MW)
    • Log of Partition Coefficient (LogP)
    • Topological Polar Surface Area (TPSA)
    • Fraction of $sp^{3}$ hybridized carbon atoms ($Fsp^{3}$)
  • Structure flexibility features. Binding sites within a targeted biomolecule (most often a protein) display by nature complex 3D geometry, with key potential sites of drug-target interactions (amino acid side chains, as an example) somewhat fixed in space. The protein, however, has a dynamic structure and even the binding pocket undergoes changes in shape. Drug-like molecules need to be sufficiently rigid to make efficient interactions with the target protein, including in most cases a high degree of selectivity over making corresponding interactions with related proteins). Perhaps less intuitive is that drug-like molecules must be flexible enough to maintain those interactions as the protein adapts its conformation. There is a "Goldilocks principle" at play, where too rigid or too flexible are each undesired extremes. Here, we chose the following two properties to account for the flexibility aspect.
    • Number of Rotatable Bonds (#rot)
    • Whether the molecule contains a macrocycle (defined as an 8-membered ring or larger)
  • Pattern-based features. In practical drug discovery, there are always some key patterns and/or scaffolds that medicinal chemists would like to hold onto or get rid of. For instance, in the lead optimization phase, retaining the key moiety and desired chemical formula are rather essential. Meanwhile, avoiding chemically unstable groups, PAINS molecules baell_chemistry_2014, and molecules that would cause structure alerts could increase the chance of success in development. Therefore, we have the following three properties for fine-tuning.
    • Avoidance of undesirable chemical patterns
    • Retention of specified substructure (between 50 and 250 Da in molecular weight)
    • Chemical formula
  • Covalent warhead feature. Drugs can be broadly categorized into noncovalent and covalent drugs, depending on whether the drug reacts with its target. For example, a the electrophilic group of a covalent inhibitor might form a bond with a nucleophilic amino acid side chain of its target protein. The reactive functional group of a covalent inhibitor is called a warhead. While most drugs are non-covalent, either can be desired. To give the model the ability to generate covalent binders, we also curated common covalent warhead-related SMARTS patterns from the Enamine fragment library 61 to indicate whether our generated molecules have the capacity to covalently bind to the target or not.
    • Whether the molecule contains common covalent warhead-related SMARTS patterns, and which of these patterns appear in the molecule
Prompting options used in fine-tuning. To incorporate the properties mentioned above into the training, we used several ways of prompting to satisfy the requirement from target uses.
For numerical properties, including all physiochemical properties and #rot, we prompted Llama by providing a specific range that the training molecules falls into for that specific category. All the cutoff values used for ranges are either commonly used standards in drug discovery or derived from the training distribution. Besides the range guidance, we added the prompt that tell Llama exactly how many #HBDs and #HBAs are contained in the training data that gives its ability to do more nuanced generation. If a property falls into multiple valid ranges—for instance, 4 H-bond donors is less than or equal to all of 4, 5 and 7—we select one of the ranges at random to place in the prompt (if the properly is picked to be placed into the prompt). It is important that the list of ranges for a property span all possible molecules, otherwise a prompt which omits information biases the model toward producing molecules with values of that property outside of the list of ranges. If we never included information in the prompt about when a molecule had over 7 H-bond donors, but did sometimes include information about the number of H-bond donors when they number under or equal to 7, then omitting the information would give the model a hint that the number of H-bond donors is more likely to be over 7. Doing this during training would bias results during inference. This is the same reason we specify undesirable properties such as the presence of bad SMARTS strings in the prompt sometimes. If the random number generator decided that a prompt should contain a substructure but the SMILES in question did not have any BRICS substructures, we added "no BRICS substructure" to the list of properties in lieu of a substructure.
For other categorical properties, we used a combination of RDKit modules, SMILES strings, and SMILES arbitrary target specification (SMARTS) strings to recognize if certain properties or chemical patterns are present in the training input. Unlike the objective of containing the scaffold exactly, chemical pattern avoidance and covalent warhead recognition required matching of more general substructures and/or certain functional groups. Here, we used SMARTS strings as our representation because of its ability of matching chemical patterns. More details about the specific SMARTS patterns used are shown later in this section.
Below is a detailed list of possible components of that could appear in a training prompt.
  • $N$ H-bond donors, $N=\leq 3,\leq 4,\leq 5,\leq 7,>7$
  • $N$ H-bond acceptors, $N=\leq 3,\leq 4,\leq 5,\leq 10,\leq 15,>15$
  • $N$ Molecular weight, $N=\leq 300,\leq 400,\leq 500,\leq 600,>600$
  • $N$ LogP, $N=\leq 3,\leq 4,\leq 5,\leq 6,>6$
  • $N$ Rotatable bonds, $N=\leq 7,\leq 10,>10$
  • $N$ Fraction $sp^{3}$, $N=<0.4,>0.4,>0.5,>0.6$
  • $N$ TPSA, $N=\leq 90,\leq 140,\leq 200,>200$
  • a macrocycle, no macrocycles
  • has bad SMARTS, lacks bad SMARTS
  • has covalent warheads, lacks covalent warheads
  • substructure of *a_smiles_string*
  • a chemical formula of *formula*
SMART patterns used to identify bad chemical groups. Li et al. pointed out a list of bad chemical patterns that exists in ChEMBL database, which will negatively affect compound generation21. In this work, we used the same list of SMARTS patterns as their work to avoid bad patterns, including cyclopentadiene, cyclopentadiene ylidenes, aromaticity-breaking tautomers, antiaromatic system, unstable halogen-heteroatom bonds, unstable fused rings, allenic system, thiazyl linkages, and peroxide bonds. In Table S1, we also present the frequency of sampling undesirable chemical groups in ChEMBL and across different generative models.
  • [C^2]1=[C^2]-[C^2]=[C^2]~[C;!d4]~[C;!^2;d2]1
  • [C^2]1~[C^2]~[C^2]~[C^2]~[C;!^2;d2]~[N]1
  • [#6^2]1~[#6^2]~[#6^3;!d4]~[#6^2]2~[#6^2]~
    [#6^2]~[#6^2]~[#6^2](~[*])~[#6^2]~2~[#6^2]~1
  • [#6]1(=[*])[#6]=[#6][#6]=[#6]1
  • [#6]1=[#6][R{2-}]=[R{2-}]1
  • [#6^2]1~[#6^2]~[#6^2]~[#6^2]~[#6^1]~[#6^1]~1
  • [#7,#8,#16]-[#9,#17,#35,#53]
  • [r3,r4]@[r5,r6]
  • [*]=[#6,#7,#8]=[*]
  • [#7,#16]=[#16]
  • [#8]-[#8]
In addition to the patterns mentioned above, we use the following SMARTS patterns to enforce our generated pyrroles to be one of the following correct forms.
  • [N^2]1~[C,N;^2](=[*])~[C,N;^2]~[C,N;^2]~[C^3]1
  • [N^2]1~[C,N;^2]~[C,N;^2](=[*])~[C,N;^2]~[C;^3]1
  • [N^2]1~[C,N;^2]~[C,N;^2]~[C,N;^2](=[*])~[C;^3]1
  • [C,N;^2](=[*])1~[N;^2]~[C,N;^2]~[C,N;^2]~[C;^3]1
  • [C,N;^2]1~[N;^2]~[C,N;^2](=[*])~[C,N;^2]~[C;^3]1
  • [C,N;^2]1~[N;^2]~[C,N;^2]~[C,N;^2](=[*])~[C;^3]1
SMART patterns used to encode common covalent warhead-related functional groups. Common covalent warheads are extracted from the Enamine Covalent Screening and Covalent Fragment Library enamine_covalent_libraries. The list of SMARTS strings is shown below.
  • sulfonyl fluorides: [#16](=[#8])(=[#8])-[#9]
  • chloroacetamides: [#8]=[#6](-[#6]-[#17])-[#7]
  • cyanoacrylamides: [#7]-[#6](=[#8])-[#6](-[#6]#[#7])=[#6]
  • epoxides: [#6]1-[#6]-[#8]-1
  • aziridines: [#6]1-[#6]-[#7]-1
  • disulfides: [#16]-[#16]
  • aldehydes: [#6](=[#8])-[#1]
  • vinyl sulfones: [#6]=[#6]-[#16](=[#8])(=[#8])-[#7]
  • boronic acids/esters: [#6]-[#5](-[#8])-[#8]
  • acrylamides: [#6]=[#6]-[#6](=[#8])-[#7]
  • cyanamides: [#6]-[#7](-[#6]#[#7])-[#6]
  • chloroFluoroAcetamides: [#7]-[#6](=[#8])-[#6](-[#9])-[#17]
  • butynamides: [#6]#[#6]-[#6](=[#8])-[#7]-[#6]
  • chloropropionamides: [#7]-[#6](=[#8])-[#6](-[#6])-[#17]
  • fluorosulfates: [#8]=[#16](=[#8])(-[#9])-[#8]
  • beta lactams: [#7]1-[#6]-[#6]-[#6]-1=[#8]
In section 3.2, we investigate SmileyLlama’s performance on the following 387 tasks, grouped into the families for which the averages are shown in the table.
  • exactly k H-bond donors, from k=0 to k=5
  • exactly k H-bond acceptors, from k=0 to k=10
  • <= k H-bond donors, for k=3,4,5,7
  • <= k H-bond acceptors, for k=3,4,5,10,15
  • <= k Molecular weight, for k=300,400,500,600
  • <= k LogP, for k=3,4,5,6
  • <= 7, <=10, >10 Rotatable bonds
  • > 0.4, > 0.5, > 0.6, < 0.4 Fraction sp3
  • <=90, <=140, <=200 TPSA
  • a macrocycle
  • no macrocycles
  • has bad SMARTS (not shown in table but included for completeness)
  • lacks bad SMARTS
  • lacks covalent warheads
  • has covalent warheads (one for each of the 16 covalent warheads in the section above)
  • a substructure of (one of each of 320 Enamine fragments 61)
  • <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 Molecular weight, <= 5 LogP
  • <= 3 H-bond donors, <= 3 H-bond acceptors, <= 300 Molecular weight, <= 3 LogP
S2 Algorithms used for SFT and DPO
Algorithm 1 Pseudocode for generating the SFT dataset
  for all SMILES_STRING $\in$ ChEMBL do
     PROMPT $\leftarrow$ "Output a SMILES string for a drug like molecule with the following properties:"
     for all (PROPERTY_FUNCTION, PROPERTY_NAME) $\in$ shuffle(PROPERTIES) do
        VALUE $\leftarrow$ PROPERTY_FUNCTION(SMILES_STRING)
        if random_coin_flip() $=$ HEADS then
           PROMPT $\leftarrow$ PROMPT $+$ PROPERTY_NAME $+$ " of " $+$ VALUE $+$ ", "
        end if
     end for
     prompts.append(PROMPT)
     completions.append(SMILES_STRING)
  end for
Algorithm 2 Pseudocode for generating the DPO dataset for improving prompt obedience
  for all (SMILES_STRING, CONDITIONS) $\in$ RESPONSES do
     if MEETS_CONDITIONS(SMILES_STRING, CONDITIONS) $=$ True then
        winners.append(SMILES_STRING)
     else
        losers.append(SMILES_STRING)
     end if
  end for
  pairs $\leftarrow$ randomly_pair(winners, losers)
  assert: $|\text{pairs}|=\min\{|\text{winners}|,|\text{losers}|\}$
Algorithm 3 Pseudocode for generating the DPO dataset for Mpro optimization every iteration
  PROMPT $\leftarrow$ "Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO"
[…]
     SMILES_LIST $\leftarrow$ SmileyLlama.inference(PROMPT, num_responses)
     SMILES_LIST $\leftarrow$ SMILES_LIST.remove_redundancies()
  end while
  for SMILES $\in$ SMILES_LIST do
     iMiner.compute_score(SMILES)
  end for
  for i $=$ 1 to 8 do
     for SMILES $\in$ SMILES_LIST do
        while True do
           RANDOM_OTHER_SMILES $\leftarrow$ pick_random(SMILES_LIST)
           if iMiner.retrieve_score(SMILES) $>$ iMiner.retrieve_score(RANDOM_OTHER_SMILES) then
              DATASET.append((PROMPT, SMILES, RANDOM_OTHER_SMILES))
              break
           else if iMiner.retrieve_score(SMILES) $<$ iMiner.retrieve_score(RANDOM_OTHER_SMILES) then
              DATASET.append((PROMPT, RANDOM_OTHER_SMILES, SMILES))
              break
           end if
        end while
     end for
  end for
S3 SmileyLlama Outside of Chemical Language Modeling
While SFT clearly alters the model, one might wonder whether SmileyLlama retains LLM abilities unrelated to exploring chemical space. We find that SmileyLlama inherits some ability to process natural language, however, SmileyLlama will sometimes respond to chemistry-related queries with a SMILES string rather than a sensible, natural-language answer. Below are four sample responses from SmileyLlama to four prompts. Each was generated with greedy decoding, which is equivalent to sampling at a temperature of 0. All responses were generated using the Alpaca instruction format, with a system prompt/instruction of “You are a helpful assistant”. As shown below, SmileyLlama is capable of writing working python code for simple functions and answering some questions in natural language. This demonstrates that, while SmileyLlama was designed to be a CLM, it still inherits some of Llama’s natural language processing ability. However, this is not perfect, as SmileyLlama will sometimes respond to chemistry-related queries with a SMILES string, rather than the correct answer.
Input: Write me some python code which calculates the n’th Fibonacci number
Output:
def fibonacci(n):
if n <= 0:
return "Input should be a positive integer"
elif n == 1:
return 0
elif n == 2:
return 1
else:
a, b = 0, 1
for _ in range(2, n):
a, b = b, a + b
return b
Input: What is the molecular weight of hydrazine?
Output: N1N=C(C)C(C1)C
(Note: This is the SMILES representation of 3,4-dimethyl 2-pyrazoline, rather than hydrazine)
Input: What is a Diels-Alder reaction in organic chemistry?
Output: A Diels-Alder reaction is a [4+2] cycloaddition between a diene and a dienophile.
Input: Who was Christopher A. Lipinski in the field of chemistry? What was his "Rule"?.
Output: Christopher A. Lipinski was a chemist who developed the "Lipinski’s Rule of Five" which states that for a drug to be orally bioavailable, it should have the following properties:
1) C(C)Cn1c2c(cccc2)c(c1)C(Nc1ccccc1)=O
2) A molecular weight of 500 or less
3) No more than 5 H-bond donors
4) No more than 10 H-bond acceptors
Furthermore, we used the Language Model Evaluation Harness to test SmileyLlama’s performance on the MMLU GPQA, Math-Hard, and MMLU-Pro benchmarks 74, 75, 76, 77, 78. The results in Table S2 and Figure S3 show that SmileyLlama generally performs worse on most tasks compared to Llama, especially difficult tasks as assessed in the latter three benchmarks. For practical purposes, the SFT procedure for creating SmileyLlama should generally be thought of as converting Llama into a CLM, rather than simply adding an additional functionality to Llama.
S4 Supporting Tables
Table S1: Frequency of generating or sampling undesirable chemical structures in ChEMBL, SmileyLlama, and other SOTA methods. Here, we randomly sample 10,000 molecules from the ChEMBL dataset, SmileyLlama and our retrained models to compare their statistics of five different categories of undesirable chemical patterns.
Model ChEMBL SmileyLlama LSTM GPT S4
Unsaturated Benzene 0.55% 0.1 % 0.04% 0.04% 0.03%
Unsaturated Naphthalene 0.02% 0.01% 0.02% 0.01% 0.01%
Wrong Pyrrole 0.29% 0.02% 0.02% 0.03% 0.04%
Cyclopentadiene Ylidene 0.96% 0.1 % 0.04% 0.06% 0.05%
Benzyne $0$ 0.0% 0.0% 0.0%
Table S2: SMILES strings and scores for selected generated molecules. All Synthetic Accessibility (SA) Scores were calculated using the method by Ertl and Schuffenhauer ertlEstimationSyntheticAccessibility2009.
SMILES string Prompt Vina score SA Score
c1(-c2cc(-c3cc(cc(c3)F)F)nc(n2)N2C[C@H](O[C@H](C2)C)C)cc2c(nc(nc2N)N)cc1 SARS2Pro -9.6 3.44
c1(nc2cc(ccc2c(n1)Nc1cc2c(cc1F)nc(nc2N)N)F)Nc1cc2c(cc1F)[nH]nc2 SARS2Pro -9.9 2.95
c1(cc(-c2cc3c(nc(nc3Nc3cc4c(cc3F)OCCO4)N)cc2)cc2c1[nH]nc2)F SARS2Pro -10.4 2.82
c1(-c2cc3c(nc(nc3Nc3cc(ccc3)C(N3C[C@H](O[C@H](C)C3)C)=O)N)cc2)cc2c([nH]nc2)cc1 SARS2Pro -10.0 3.36
C1N(c2cc(-c3cc4c(nc(nc4N)N)cc3)nc(n2)Nc2cc3c(cc2F)occ3)CCOC1 SARS2Pro+Ro5 -9.6 2.86
c1cc2cc(-c3nc4c(ncnc4Nc4cc5c(cc4F)OCCO5)cc3)ccc2nc1N SARS2Pro+Ro5 -10.2 2.68
c1cc(-c2cc3c(nc(nc3Nc3cc4c(cc3F)OCCO4)N)cc2)cc2c1[nH]nc2C SARS2Pro+Ro5 -10.2 2.72
C1(N(c2cc(-c3nc4c(cccc4)c(n3)Nc3cc4c([nH]nc4)cc3)cc(c2)F)C(=O)CC1)=O SARS2Pro+Ro5 -10.1 2.64
Table S3: Performance of Meta-Llama-3.1-8B-Instruct versus SmileyLlama in a variety of LLM benchmarks. We report the mean score, assessed with the exact_match scoring method.
Model Llama-3.1-8B-Instruct SmileyLlama
MMLU (0-shot) 0.68 0.65
GPQA (chain of thought) 0.30 0.038
Math-Hard 0.24 0.039
MMLU-Pro 0.47 0.063
S5 Supporting Figures
Figure S1: Guacamol benchmark results as a function of temperature. We used n=3 runs to assess how validity, uniqueness and novelty depends on temperature, a parameter which affects the randomness of outputs generated during inference. Error bars are defined by the standard error. Temperature can be increased during inference for more random molecules, sacrificing validity for diversity and novelty. We generally use a temperature of T=1.0 during this study.
Figure S2: Validity, uniqueness and novelty over optimization for SARS2-Mpro inhibition We plot the validity (fraction of molecules which are valid), uniqueness (number of distinct molecules) and novelty (fraction of molecules which have not been generated before) of molecules sampled from the generative models of iMiner with 2000 molecules per epoch, and for SmileyLlama with DPO (SL+DPO) with 2000 and 1000 molecules per epoch. Note that novelty is not defined with respect to ChEMBL; here it’s a measure of how good the model is at generating molecules that it has not generated in a previous iteration.
Figure S3: SmileyLlama de novo generated molecules in the active site of SARS2 main protease. Surface rendering of the SmileyLlama generated molecules in the SARS2 Mpro canonical binding pocket. Generated by SmileyLlama after optimization with the SARS2PRO and SARS2Pro+Ro5 prompts. Table S3 provides their SMILES string and docking scores.
Figure S4: A visualization of how the SmileyLlama SFT procedure affects zero-shot Massive Multitask Language Understanding (MMLU) score75. Generally, MMLU scores appear to decrease with fine-tuning. One might expect that fine-tuning which improves the ability to generate molecules may translate to improved performance on chemistry, but this was found to be the opposite of the case; chemistry-specific sections of the MMLU saw among the worst degradation in score after SFT. Anecdotally, SmileyLlama will often respond to a chemistry-related query with a SMILES string. The sharpest decrease in performance was in the "Moral Scenarios" section. This may be due to much of the "moral knowledge" coming during post-training. The "moral knowledge" section of the MMLU had the largest difference in score between Between the base and instruct Llama models; while SmileyLlama scores about 41% compared to Llama-3.1-8B Instruct’s 57%, the base model Llama-3.1-8B score centers around 28%. On the other hand, there are some tasks, such as "Machine Learning", where performance appears to improve, although the standard error bars for each of these tasks include the mean so we cannot conclusively say that the SFT procedure used for SmileyLlama improved Llama’s abilities outside of SMILES string generation.