### v4

# SmileyLlama: Modifying Large Language Models for Directed Chemical Space Exploration

Abstract
Here we show that a general-purpose large language model (LLM) chatbot, Llama-3.1-8B-Instruct, can be transformed via supervised fine-tuning of engineered prompts into a chemical language model (CLM), SmileyLlama, for molecule generation. We benchmark SmileyLlama by comparing it to CLMs trained from scratch on large amounts of ChEMBL data for their ability to generate valid and novel drug-like molecules. We also use direct preference optimization to both improve SmileyLlama’s adherence to a prompt and to generate molecules within the iMiner reinforcement learning framework to predict new drug molecules with optimized 3D conformations and high binding affinity to drug targets, illustrated with the SARS-Cov-2 Main Protease. This overall framework allows a LLM to speak directly as a CLM which can generate molecules with user-specified properties, rather than acting only as a chatbot with knowledge of chemistry or as a helpful virtual assistant. While our dataset and analyses are geared toward drug discovery, this general procedure can be extended to other chemical applications such as chemical synthesis.
1Kenneth S. Pitzer Theory Center and Department of Chemistry,
University of California, Berkeley, CA, 94720 USA
2Department of Computer Science, University of Wisconsin–Madison, Madison, WI, 53706
3Department of Molecular Medicine, The Herbert Wertheim UF Scripps Institute for Biomedical Innovation and Technology, 130 Scripps Way, Jupiter, FL, 33458 USA
4Departments of Bioengineering and Chemical and Biomolecular Engineering,
University of California, Berkeley, CA, 94720 USA
5Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA
$\dagger$ thg@berkeley.edu
1 Introduction
Chemical Language Models (CLMs)grisoni_chemical_2023 trained on string representations such as Simplified Molecular-Input Line-Entry System (SMILES) smiles and SELF-referencIng Embedded Strings (SELFIES) selfies have emerged as a useful tool for de novo generation of molecules, best exemplified by molecules relevant to pharmaceutical applications and drug discoverymol-gen-bg0, mol-gen-bg1, mol-gen-bg2, mol-gen-bg3, mol-gen-bg4, mol-gen-bg5. Nearly all CLMs for molecular generation have been trained from scratch on large quantities of data such as ChEMBLgaulton_ChEMBL_2012 and ZINCTingle2023 and using different model architectures, including variational autoencodersVAE2018, recurrent neural networks (RNNs) long short term memory (LSTM) cells lstm, mol-gen-rnn, generative pre-trained transformers (GPTs)gpt, mol-gen-gpt0 and structured state space sequence (S4) modelss4, Ozcelik2024s4. In addition, CLMs have achieved advances in chemical generation tasks through further downstream optimization of molecules with additional training or different model frameworks.mol-gen-bg3, mol-gen-bg4, mol-gen-gpt1, mol-opt1, iminer A recent such example is the iMiner algorithm that operates on SELFIES and uses deep reinforcement learning (RL) with real-time 3D docking to generate molecules with high potential for inhibiting protein targets.iminer
Language models are statistical models of probability distributions of units of language, and can be adapted to generate meaningful text by sampling from these distributions.880083 The most recent advancement of language models have resulted from the training of scaled-up transformers vaswani2023attentionneed on massive amounts of data, resulting in the creation of Large Language Models (LLMs)radfordllms such as the proprietary GPT-4openai_gpt-4_2024 and the open-weight Llama.dubey_llama_2024. While LLMs were originally created for applications in natural language processing, in recent years there have also been efforts to adapt LLMs to chemistry. Some studies have accessed frontier LLMs for the purpose of assisting research in the form of virtual lab members, or even performing research autonomously.Zheng2023, boiko_autonomous_2023, m_bran_augmenting_2024 Other studies have focused on training LLM architectures on chemical data which can translate between natural and chemical languagesmolT5, pei2023BioT5 such as finding a molecule that is a close match to a ground truth molecule from which the description is assessed by fingerprint tanimoto similarity.yu2024llasmol Beyond chemical dictionary lookups, LLMs have also been used to perform mutation and crossover for an evolutionary algorithm to explore chemical space.molleo There have also been studies of LLMs to modify SMILES strings to change the properties of the molecules that they represent.guevorguian_small_2024, mol-llm-engine-1, drugllm, wu_leveraging_2024 Others have taken inspiration from LLMs to design CLMs with new capabilities, such as the ability to respond to prompts through a transformer-based architecturewu_leveraging_2024, mol-llm-engine-2. However, to our knowledge, no CLM derived from a pre-trained general purpose LLM foundation model has reached the performance exhibited by modern CLMs that are trained from scratch with chemical data.
Here, we demonstrate that an open-weight LLM, Meta-Llama-3.1-8B-Instruct (which we will refer to as “Llama” for the remainder of the paper) dubey_llama_2024, can be converted into a CLM for generative tasks in drug discovery. The fact that Llama is open-weight offers several benefits as a foundation model. In particular, this allows users to train and share adapters, to perform training and inference without needing to store potentially valuable data on a remote server, to have complete control over the hyperparameters and algorithms used for fine-tuning, and to perform interpretability analyses on the model weights. Using supervised fine-tuning (SFT) and Direct Preference Optimization (DPO)rafailov_direct_2024 of the pre-trained Llama with SMILES strings derived from ChEMBL, SmileyLlama generates drug-like molecules with desirable properties specified in a user-defined prompt with relevance to medicinal chemistry. SmileyLlama not only uses the ability of LLMs to have their behavior guided by prompts, but also their ability to serve as a foundation model which can be fine-tuned to specialize in various text-generation tasks for molecule generation that we show can match or exceed the performance of modern CLMs trained directly on chemical data. We further demonstrate that SmileyLlama greatly improves the reinforcement learning component of iMiner to more efficiently explore chemical space to create molecules optimized for 3D binding to target proteins, illustrated with the SARS-Cov-2 (SARS2) main protease (MPro).mpro_review While our dataset and subsequent analyses are created with drug discovery in mind as a downstream application, this general procedure can be extended to other chemical applications such as chemical synthesisSun2025.
2 Methods
2.1 Supervised Fine-Tuning and Direct Preference Optimization
In order to steer the outputs of the pre-trained Llama model dubey_llama_2024 for drug molecule generation we first use supervised fine-tuning (SFT), in which the weights of Llama are further optimized on generic drug-like properties (Section 3.1). We constructed our fine-tuning dataset with the SMILES strings of approximately 2 million molecules from ChEMBL Dataset v33 gaulton_ChEMBL_2012. For each molecule in our dataset, we randomly picked a number of properties of pharmaceutical interests to calculate using RDKit Landrum2016RDKit2016_09_4. These properties include ranges of hydrogen-bond (H-bond) donors and acceptors, molecular weight (MW), log of the partition coefficient (logP), number of rotatable bonds, the fraction of $sp^{3}$ carbons (i.e., the number of $sp^{3}$ carbons divided by the total number of carbon atoms in the molecule, termed $Fsp^{3}$), the presence/absence of macrocycles, the presence/absence of covalent warhead-related SMARTS patterns, the presence/absence of at least one undesirable SMARTS pattern, a SMILES string representation of a BRICS single-pass substructure and the specification of a chemical formula.
For numerical properties, we choose ranges that are relevant for medicinal chemistry. In many medicinal applications, molecules must have suitable characteristics of drugs related not just to their chemical architecture but also to relevant biological phenomena (e.g., oral absorption, metabolism, distribution). For example, molecules invented in the early stages of drug design are typically smaller in size and lipophilicity than their follow-on analogs and optimized drugs, so a user can specify that a model obeys the Rule of Three—a common rule of thumb for early small molecule drug design—when generating moleculesjhoti_rule_2013. As another example, we choose the TPSA ranges based on those which tend to be orally bioavailable or able to pass through the placenta or blood-brain-barrierveber_molecular_2002. If a drug need not meet these criteria (e.g., an injectable drug for a peripheral target may not require as specific a TPSA range) then a user interfacing with SmielyLlama should be able to adjust the range criterion or eliminate it. Further specifics of these properties and the ranges we choose to specify during training can be found in the Supplementary Information in Section 1.
After calculating and picking these properties for each SMILES string, we construct a prompt containing values of these properties, with the “correct” completion being the SMILES string that these properties were calculated from. For example, we used a prompt with a system instruction of You love and excel at generating SMILES strings of drug-like molecules and a user instruction of the form Output a SMILES string for a drug like molecule with the following properties: if properties are specified, or Output a SMILES string for a drug like molecule: if no properties are specified. We choose to create prompts that assign SmileyLlama the role of an AI which excels at producing SMILES strings due to the track record of role-promptingchenUnleashingPotentialPrompt2025; we chose this prompt format due to its balance between motivation and brevity. Each property has a 50% chance of being calculated and specified in the prompt so that the model can operate equally well during inference, whether or not any properties are specified. This trains our model to generate molecules that have properties specified (or not) in the prompt. We structure the prompts used for SFT so that, during inference, the properties of generated molecules can be specified in the prompt. This lets users avoid having to downselect the vast majority of generated molecules for having the correct characteristics—instead, users can simply prompt SmileyLlama to provide molecules with the characteristics they desire. See Figure 1 and Algorithm 1 in the Supporting Information for further elaboration of SFT training.
Figure 1: A visualization of the SFT workflow for Smiley-Llama. Given the Llama-3.1-8B-Instruct model dubey_llama_2024, we used prompt-response pairs consisting of calculated molecular properties and completed SMILES strings to fine-tune Llama on SMILES strings completions, yielding SmileyLlama. Crucially, we construct the prompt for each example using properties calculated from the correct response (a SMILES string from ChEMBLv33).
We also use Direct Preference Optimization (DPO)rafailov_direct_2024 which also updates the weights of Llama such that the language model is more likely to produce strings designated as "winners" and less likely to produce strings designated as "losers". We use DPO to further reinforce our model’s ability to robustly generate molecules with properties specified in the prompt for more specific task-oriented goals (Sections 3.2 and 3.3). Algorithmically we prompt our SFT model to generate molecules with a given property, and sample several SMILES strings and use RDKitLandrum2016RDKit2016_09_4 to assess whether they have properties in line with what the prompt requested. We then pair up molecules which correctly followed the prompt with those that don’t as winners and losers, respectively, then use a single epoch of DPO to improve the model’s results. See Algorithm 2 in SI for pseudocode of this scoring and pairing procedure.
Finally we consider DPO for generating unique and valid ligands which bind to a specific protein target, Sars2-Mpro using the iMiner framework, the details of which have been published elsewhere iminer. During our optimization process, we prompt SmileyLlama to provide molecules with a “High SARS2PRO” score. After every 2000 distinct molecules are generated per iteration, we score each molecule and randomly pair it with another. The molecule with the higher score is designated as the "winner" and the other as the "loser", and we construct a dataset with eight such pairings for each molecule and then optimize SmileyLlama using DPO with respect to this dataset. We empirically found 8 pairs to work well; more than 8 pairs tended to lead to a lack of diversity, whereas fewer than 8 pairs resulted in less sample efficient optimization. We use a scoring function of three times the docking score plus the iMiner drug-likeness functioniminer to score all molecules per iteration. Invalid molecules were assigned a score of -10. This process is described in Algorithm S3 in the Supplementary Information.
2.2 Additional Training Details
We performed both SFT and DPO on Llama using the Axolotl Package dubey_llama_2024, noauthor_axolotl-ai-cloudaxolotl_nodate. For both SFT and DPO, we use the Low-Rank Adaptation (LoRA) applied to the linear layers of the model and FlashAttention with an Adam optimizer, cross-entropy loss, and a cosine learning rate scheduler with a maximum learning rate of $2\times 10^{-4}$.hu_lora_2021, dao2022flashattention, dao2023flashattention2, kingma_adam_2017 All prompts were formatted according to the Alpaca instruction format alpaca, wang-etal-2023-self-instruct. Additional parameters for our training are a LoRA rank of 32, a LoRA alpha of 16, a LoRA dropout of 5%, and 10 warmup steps. We inherit these hyperparameters from standard practice with Axolotl. For example, LoRA hyperparameters identical to these were used in the Hermes 3 SFTHermes3TechnicalReport. For SFT, we trained for 1 epoch using a batch size of 64 on a single 4xA40 node for approximately 32 hours with a validation set $\sim$5% the size of the original, an amount which would cost approximately \$53 in October 2025 on Vast.ai. We also note that we randomized the SMILES string representation of each molecule, and we tokenized all SMILES strings with the Llama3 tokenizerdubey_llama_2024 when interfacing with SmileyLlama.
3 Results
3.1 Comparison of SmileyLlama from SFT with other CLMs
To test the generative ability of SmileyLlama compared with other existing CLMs across a range of architectures, we used the GuacaMol suite brown2019guacamol to benchmark the validity, uniqueness, and novelty of the molecules as shown in Table 1. Additionally, KL divergence and Frechet Chemnet Distance preuer2018frechet based on the GuacaMol definition (FCDGuac) are used to analyze the distributional shifts from the ChEMBL training data for drug-like molecules more specifically. The Fréchet ChemNet Distance has been introduced as a metric for chemical and biological relevance for drug discoverypreuer2018frechet, and we note that FCDGuac=exp(-0.2*FCD) compresses the FCD distances themselves for the convenience of defining a 0 to 1 value like the other Guacamol metricsbrown2019guacamol. Hence the scores should actually be considered from a log perspective such that an FCD below 5 (i.e. FCDGuac =0.37) is in strong agreement with distributions of drug-like properties. Only when the FCDGuac score drops by one to two orders of magnitude are they considered to be poor FCD scores. We recommend in the future that others report the more straightforward FCD distances themselves to avoid this interpretative confusion.
[…]
Benchmark Validity Uniqueness Novelty KL div FCDGuac
GraphMCTSIwata2023 1.000 1.000 0.994 0.522 0.015
VGAE-MCTSIwata2023 1.000 1.000 1.000 0.659 0.009
AAEIwata2023 0.822 1.000 0.998 0.886 0.529
LSTMOzcelik2024s4 0.983 0.999 0.848 0.993 0.901
GPTOzcelik2024s4 0.915 1.000 0.978 0.977 0.826
S4 Ozcelik2024s4 0.971 0.997 0.961 0.994 0.853
Llama zero-shot 0.688 0.457 0.635 0.736 0.002
Llama twenty-shot 0.465 0.999 0.949 0.913 0.079
SmileyLlama 0.958 1.000 0.987 0.967 0.686
We first analyze the ability of Llama to produce molecules, relying only on its pre-trained knowledge (zero-shot), or by providing it with one or more examples from the ChEMBL database in the formulated prompt (Table 1 and Supplementary Table S1). We find that without SFT or examples provided in a prompt the LLM is unable to produce high percentages of valid SMILES strings compared to other state-of-the-art CLMs, and is generally poor even with variations in hyperparameters such as temperature ($T$). Interestingly, validity is lower when 20 examples are provided in the prompt (twenty-shot) than it is when no examples are in the prompt (zero-shot). We speculate that Llama zero-shot has had some exposure to the SMILES syntax to be able to generate valid strings, but it has no intrinsic ability to generalize and it falls back to repeating the memorized SMILES and resulting in low uniqueness. When several examples are given in the prompt, this biases Llama away from the known SMILES strings it can produce, but the few examples means that its grasp on the allowed mutable structure of SMILES strings is poor and thus less valid. However, because these prompts are so diverse, Llama’s 20-shot uniqueness is very high.
In Table 1 it is seen that SFT significantly improves SmileyLlama’s ability to generate drug-like molecules as well as the CLMs. In addition, we experiment with the format of the SmileyLlama prompt, performing SFT on Llama with a less anthropomorphic user prompt and a blank template as an ablation study. We find that changing this prompt format does not significantly affect the guacamol benchmarks. More information on this prompt format, as well as that for the Llama models, can be found in Supplementary Information section 1.3 and Table S1. To show the generality of the LLM-SFT approach, we also fine-tune Llama-3.2-3B, Llama-3.2-1B, and Qwen-2.5-7Bqwen25TechnicalReport2025 using the same SFT workflow (including identical hyperparameters) that we developed for SmileyLlama, yielding the models we name SmileyLlama-1B, SmileyLlama-3B, and SmileyQwen2.5-7B, respectively. Supplementary Table S1 finds that the Guacamol benchmark results did not change substantially between SmileyLlama and SmileyQwen2.5-7B. We also find through inspecting the results of SmileyLlama-1B and SmileyLlama-3B, that the validity increases with parameter count, though novelty, uniqueness and the match between the training distribution and the distribution of generated molecules stays roughly unchanged. While our later analyses are restricted to SmileyLlama, these results suggest that a more capable, future LLMs could also be converted into a CLM using an identical or similar SFT method.
Figure 2: Distribution comparisons for different properties of the generated molecules from SmileyLlama (blue) with molecules from the training dataset from ChEMBL (gold). (a) UMAP visualization of a random selection of 10,000 ChEMBL molecules and 10,000 SmileyLlama-generated molecules, using 15 neighbors and a minimum distance of 0.1; these are normal values in chemical space visualizationorlovHighDimensionsHuman2025. (b) The molecular properties considered are fraction of $sp^{3}$ hybridized carbons and heteroatoms, number of heavy atoms, number of H-bond donors and acceptors, number of aliphatic and aromatic rings and the maximum ring size, number of rotatable bonds, quantitative estimate of drug-likelihood (QED) valuebickerton2012quantifying, MW, approximate log partition coefficient between octanol and water (ALOGP)wildman1999prediction, polarizable surface area (PSA) and topological PSA Prasanna2009, and the number of structural alerts brenk2008lessons. All benchmarks were at a temperature T=T=1.0 and a maximum of 256 new tokens.
Figure 2 shows that SmileyLlama generates very good agreement with ChEMBL quantities across a diverse property set. The UMAP visualization in Figure 2a, a popular visualization tool used in drug discovery, finds that while there are small relative differences in densities, SmileyLlama generates molecules in every well-represented region of the chemical space of ChEMBL. Like all data reduction techniques, UMAP representations are qualitative, and hence we also consider the distribution of molecular properties of interest to medicinal chemistry in Figure 2b, and KL-divergence values indicates all properties are in very good agreement between the SmileyLlama-generated molecules and ChEMBL molecules. Hence there are no obvious chemical regions of ChEMBL that are inaccessible to SmileyLlama. Furthermore, it has recently come to our attention that certain undesirable molecular scaffolds might be oversampled in CLM generative tasks.iminer Supplementary Table S1 shows that although small percentages of these unviable chemical structures are present in the ChEMBL training data itself, all CLMs do not oversample these undesirable molecular scaffolds. Finally, while training was conducted at $T=1.0$, exploration of temperature used at inference on the Guacamol benchmark (Supplementary Fig. 1) suggests that this temperature is adequate for all tests described in the Results.
3.2 Property Specification
In Table 2 we show the average percentage of valid, distinct SMILES strings generated for a complete panel of families of tasks. This benchmark is distinct from other conditional molecule generation benchmarksmol-gen-gpt0, yu2024llasmol in that we are testing SmileyLlama’s ability to robustly generate molecules with properties in value ranges rather than molecules with a property of a specific value. This is for the reason of what is of interest to medicinal chemistry; many of the "rules of thumb", such as counting the number of Lipinski violations or number of hydrogen bond donors and acceptors, are based on value ranges during chemical exploration. Additionally, LLMs tend to struggle with numbers which have many degrees of precision and must be split into several tokensschwartz-etal-2024-numerologic. Hence we did not represent this category in the prompt during training.
We compare SmileyLlama, both before and after DPO, with a model resulting from an ablation study on the efficacy of prompting. We study this by removing all indications of molecular properties from all of the prompts in the dataset used to train SmileyLlama; each molecule from ChEMBL is a completion to the same prompt (the prompt used for SmileyLlama when no properties of a molecule are selected). We then run SFT with exactly the same hyperparameters as SmileyLlama. Finally, we perform the same analysis of how well this model follows properties given in its prompt as we ran with SmileyLlama both before and after DPO. We find that this model performs quite poorly in comparison to SmileyLlama in this benchmark, only achieving 90+% performance on three tasks. This becomes especially pronounced when the properties are rarely found in the data, such as the presence of a macrocycle or a warhead-related SMARTS pattern. The stark contrast in performances speaks to the necessity of our prompt engineering scheme; we cannot rely purely on the knowledge of the foundation model when fine-tuning for chemical tasks.
Overall SmileyLlama model does very well on tasks on which it was trained through the engineered prompt, especially when contrasted with the model resulting from the "prompt ablation" experiment. Although some of the properties were present in training, such as the Lipinski rule-of-5 and the presence of a macrocycle, both were more uncommon in the data and hence performance in these categories are less good. Similarly while there were many warhead-related SMARTS and Enamine substructures that were specified in the SFT training procedure, some were more common than others and performance is thus more middling as also seen in Table 2. Additionally, some tasks did not occur at all in the SFT training data, and yet are still parsible as valid SMILES. Finally, and as expected, SmileyLlama does poorly on tasks involving exact numerical specifications. We note that one has a choice to use SmileyLlama using lower temperatures at inference that can improve the SFT predictions as also seen in Table 2.
Table 2: Percentage of valid, distinct generated molecules over a more complete panel of tasks using SmileyLlama. We performed a panel of 380 tasks (see Supplementary Information for complete definition) to assess how well SmileyLlama follows directions in a prompt. For each task, we generate 1000 molecules from a prompt requesting some property, and score the result based on the proportion of molecules which are valid, distinct, and satisfy the properties requested in the prompt. Finally, we collect some tasks into families (such as those with different range values) and average their scores to produce the results shown in the table below. All benchmarks were at a temperature T=T=1.0 and a maximum of 128 new tokens to increase computational efficiency. We also note that running SFT for two epochs rather than one did not seem to greatly affect the performance on these benchmarks.
SFT DPO Prompt Ablation
Present in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0 T=0.7 T=1.0
$\leq$ k H-bond donors 96.6% 94.4% 99.2% 98.1 % 93.8% 91.8%
$\leq$ k H-bond acceptors 96.3% 90.7% 98.3% 98.4 % 76.6% 65.4%
$\leq$ k Molecular weight 90.1% 84.3% 97.8% 98.0 % 62.4% 58.8%
$\leq$ k ClogP 89.2% 85.4% 96.8% 97.7 % 67.9% 64.9%
Rotatable bonds in range 90.0% 86.7% 94.5% 94.0 % 62.2% 58.5%
Fraction $sp^{3}$ in range 87.6% 85.0% 96.3% 95.9 % 30.7% 31.8%
TPSA in range 95.9% 91.1% 98.8% 98.4 % 89.2% 83.1%
No bad SMARTS 92.5% 89.6% 94.9% 94.3 % 87.3% 85.6%
Absence of macrocycle 98.4% 95.3% 98.4% 97.3 % 97.1% 94.8%
Absence of warhead-related SMARTS 96.4% 94.2% 97.3% 95.4 % 94.6% 92.9%
Lipinski rule-of-five 89.0% 81.7% 98.7% 97.7 % 73.6% 64.0%
Presence of warhead-related SMARTS 58.3% 51.0% 74.9% 73.0 % 0.2% 0.4%
Presence of Enamine substructures 51.0% 51.5% 67.7% 70.1 % 2.7% 1.9%
Presence of macrocycle 38.8% 44.7% 58.0% 57.8 % 2.1% 2.1%
Absent in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0 T=0.7 T=1.0
Rule-of-three 77.5% 62.0% 84.8% 94.5% 7.5% 5.5%
Exactly k H-bond donors 21.4% 19.7% 30.7% 30.4% 17.4% 16.8%
Exactly k H-bond acceptors 19.9% 14.4% 27.0% 31.9% 9.2% 8.8%
SmileyLlama also performs well in driving not just singular molecular design tasks, but on compound tasks such as generating molecules similar to existing leads, i.e. “scaffold hopping” R-group modification, and/or structure-based design to grow molecules from ligand fragments. Figure 3a is an example of SmileyLlama model’s ability to generate molecules from all 320 substructures in the Enamine databaseenamine_essential_libraries that follow the Lipinski rule-of-fiveLIPINSKI20013, which encompasses most of the molecular properties with ranges listed in Table 2.
While the SFT SmileyLlama model typically does well on tasks on which it was trained through the engineered prompt, and can still perform adequately when queried with prompts different from ones on which it was trained (Table 2), we can further optimize SmileyLlama for specific tasks using DPO. DPO’s most popular application has been in improving the responses of LLM-derived chatbots, such as the instruct-tuned Llama3 models, but it has also found use in improving the outputs of CLMs park_preference_2023 and avoiding the need to separately train a reward model rafailov_direct_2024. Here the relevance of DPO provides a way to further optimize the model by pairing desirable responses with undesirable responses. The model’s weights are then updated to be more likely to produce the ‘winner’ of the pairing and less likely to produce the ‘loser’ of the pairing. We generated our dataset by simply pairing up unsuccessful attempts at generating structures with successful attempts randomly for each task in Table 2. We concatenate these pairings into a single dataset and optimize for 1 epoch using DPO.
SmileyLlama optimized with DPO significantly improved adherence to the prompt across nearly all tasks as seen in Table 2 and Figure 3b. Note that while DPO does cause the model to more robustly obey the rules in the prompt, it also shifts and narrows the property distribution compared to the training set as seen in Figure 3b and appears to be largely insensitive to temperature. SmileyLlama without DPO, on the other hand, occasionally does not obey the prompt but more faithfully reproduces the distribution of properties found in a filtered ChEMBL that satisfy Lipinski’s rules. In the context of drug discovery, SFT is primarily beneficial for early exploration of chemical space, whereas DPO is a type of constraint optimization that limits generated molecules to desired sub-classes specified by the user.
Figure 3: Conditional generation with SmileyLlama for fragment growth and before and after DPO compared to ChEMBL. (a) Example molecules generated by growing from one of the Enamine substructures and to satisfy Lipinski’s Rule of 5 using the prompt Output a SMILES string for a drug like molecule with the following properties: a substructure of O=C(O)c1ccc(C(F)(F)F)cc1, <= 500 MW, <=5 logP, <= 5 H-bond donors, <= 10 H-bond acceptors. (b) distribution of four properties satisfying Lipinski’s rule of five comparing ChEMBL molecules (orange) with molecules generated by SmileyLlama (blue) with the prompt Output a SMILES string for a drug like molecule with the following properties: <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 MW, <= 5 logP, compared to 1000 molecules generated by SmileyLlama with the same prompt after DPO (gray). MW and LogP distributions were estimated using a gaussian kernel density estimator (KDE).KernelDensityEstimators1992 All results generated 1000 molecules at a temperature T=T=1.0 and a maximum of 128 new tokens.
3.3 Optimization of Binding Affinity to Protein Active Sites with SmileyLlama/iMiner
The tests performed in Sections 3.1 and 3.2 do not take advantage of the 3-dimensional structural information of a putative drug nor its shape and molecular compatibility with a target protein active site. Hence we use SmileyLlama to generate unique and valid ligands which undergo further optimization for binding to a specific protein target when embedded in the iMiner frameworkiminer. iMiner generates novel inhibitor molecules for target proteins by combining deep reinforcement learningolivecrona2017molecular, popova2018deep with real-time 3D molecular docking using AutoDock VinaTrott2010, thereby simultaneously creating chemical novelty while constraining molecules for shape and molecular compatibility with target active sites. More specifically, the iMiner generative model uses an Average Stochastic Gradient Descent Weigh-Dropped Long Short-Term Memory (AWD-LSTM)awd-lstm recurrent neural network that predicts probability of string tokens to concatenate on a molecular string representation until a complete molecule is generated. In the subsequent RL stage, 2000 molecules in each epoch (typically $\sim$ 50 epochs) are sent to AutoDock simulations in parallel, and the docking scores are circled back to the RL model to adjust its parameters so that molecules generated in the next iteration will have better scores while retaining drug-likeness.
Given the ascendancy of attention mechanisms and transformer architectures, such as those inherent in SmileyLlama, we replaced the generative AWD-LSTM component of iMiner with SmileyLlama, and replaced the iMiner optimization algorithm, Proximal Policy Optimization (PPO)ppo-schulman with DPO. This is largely due to the high memory requirement of PPO; since DPO does not need to fit a reward model to the predicted and realized rewards it requires far less memory than PPO. This becomes even more true when tuning a large model compared to a small one, since generally the reward model trained is the same size as the language model used to generate strings.
To validate the effectiveness of SmileyLlama in the iMiner context, we generate inhibitor molecules for the SARS-CoV-2 Main Protease(MPro) as an example. One route to development of antiviral treatments to Covid-19 is through inhibition of MPro, an enzyme whose function is essential to the SARS-CoV-2 lifecycle.jin2020structure We chose this system not only due to its current relevance, but also because of the readily available experimental 3D structuresjin2020structure, zhang2020crystal which provides the necessary information needed for structure-based ligand design using SmileyLlama and comparisons to other successful inhibitor molecules for this target.
For the unconditional de novo generation case, SmileyLlama learns the user prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO which pertains to minimizing the AutoDock Vina score while maximizing the drug likeliness score ($S_{DL}$) of the original iMiner reward functioniminer. Figure 4a compares the docking scores of the original iMiner algorithm against SmileyLlama as a function of epoch number and with number of generated molecules per iteration. We notice first an improved data efficiency compared to iMiner, in which SmileyLlama requires roughly one-quarter of the epochs to reach a similar level of improved docking score. Due to iMiner’s use of SELFIES as its molecular representation, its validity is unparalleled, however, in later epochs, diversity crashes which explains the sharpening peaks in later iterations (Figure 4) and quantified further in Supplementary Figure S2 against the Guacamol benchmarks. This simply reflects convergence in the docking score, i.e. there are fewer novel molecules as docking score reaches the highest values. By contrast, SmileyLlama maintains more diversity while greatly improving the docking score to iMiner (Figure 4a) with some degradation in validity compared to iMiner. At 15 epochs, SmileyLlama has 93% validity, compared to iMiner’s 100% validity (Supplementary Figure S4) likely due to the fact that SmileyLlama represents molecules with SMILES strings while iMiner uses SELFIES.
Figure 4: Comparison of the shift in docking score distributions for iMiner compared to SmileyLlama over optimization epochs as illustrated for SARS2-MPro. (a) For iMiner, in later epochs diversity crashes which explains the sharpening peaks in later iterations. SmileyLlama with DPO (SL+DPO) enforces diversity throughout the optimizations (Algorithm S3), which accounts for the broad peaks, and shows superior data efficiency relative to iMiner. (b) We compare two different user prompts: Sars2Pro and Sars2Pro+Ro5. All results were generated with 2000 valid SMILES at a temperature of T=T=1.0 and a maximum of 128 new tokens.
Figure 4b shows the property distributions of the final optimized set of novel molecules from SmileyLlama from the prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO. While the property distributions are satisfactory for parts of the Lipinski Rule of 5, such as number of hydrogen bond donors and acceptors, the MW and logP results are not conforming to drug-like values. This indicates some inadequacy of the iMiner reward function, i.e. a non-optimal weighting of docking scores compared to $S_{DL}$. While a standard CLM such as iMiner would require a reweighting and/or new terms in the loss/reward function, other hyperparameter tuning, and expensive retraining, a unique advantage of SmileyLlama is that the distribution of generated molecules’ properties can be shifted using nothing more than prompt engineering, with no retraining required. Figure 4b shows that combining prompts such as Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO, <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 molecular weight, <= 5 logP, which we will refer to as the High SARS2Pro+Ro5 prompt improves properties such as MW and logP and drug-likeness scores significantly with some expected loss in high docking scores since smaller molecules make fewer intermolecular interactions.
[…]
Figure 5: SmileyLlama de novo generated molecules in the active site of SARS2 main protease. Surface rendering of the SmileyLlama generated molecules in the SARS2 Mpro canonical binding pocket. Generated by SmileyLlama after optimization with (a) the SARS2PRO prompt. (b) and (c) the SARS2Pro+Ro5 prompt. Supplementary Table S2 provides their SMILES string and docking scores, and Supplementary Figure S3 shows their docking pose, for some of the highest scoring ligands.
3.4 SmileyLlama Outside of Chemical Language Modeling
While SFT and DPO clearly alters Llama in the creation of SmileyLlama, one might wonder whether SmileyLlama still retains the full LLM knowledge of the Llama-3.1-8B model from which it derives. We find that SmileyLlama can still converse in English if it is prompted to do so, and some sample conversations are included in Section 3 in the Supplementary Information. However, we notice that its responses are often shorter than those of Llama-3.1-8B and can contain SMILES string non-sequiturs, which seem to be more likely when dealing with prompts relating to chemistry. As a more quantitative measure of its residual abilities, we evaluate performance of SmileyLlama on the Language Model Evaluation Harness containing the MMLU GPQA, Math-Hard, and MMLU-Pro benchmarkseval-harness, mmlu, wang2024mmlupro, rein2023gpqa, hendrycks2021math. Supplementary Table S3 and Figure S4 show that SmileyLlama generally performs worse on moral scenarios, and amusingly SmileyLlama also performs worse on subjects dealing with chemistry than Llama. This in part is likely due to the tendency for SmileyLlama to complete prompts relating to chemistry with a SMILES string as noted above. In addition, accuracy errors in the MMLU tests have also been noted recentlygema2025mmlu, and thus SmileyLlama’s degraded performance in chemistry may also be in part an artifact of poor testing benchmarks. Overall this result is somewhat encouraging, since it implies the possibility that LLM-derived CLMs can inherit and take advantage of the natural language processing ability of their foundation model. SmileyLlama already does this— we can steer the properties of the molecules it generates and the chemical space it explores using natural language prompts while still retaining some ability to process non-chemical natural language. However, more work will be required to view SmileyLlama as an additional capability of an LLM, as opposed to its current and primary benefit as a CLM.
Discussion and Conclusion
We have investigated the adaptation of LLMs to molecular chemistry problems that have been largely in the domain of CLMs. We find that the key advantage that LLMs have when performing as CLMs comes from prompting: information about the region of interest in chemical space can be specified in the prompt, letting the model generate chemical structures in this region. Furthermore, we showed that training an LLM on a variety of prompts allows for good performance in multiple CLM tasks, and found evidence to suggest that the LLM models can correctly generalize to scenarios with new prompts that leverage the abilities developed when trained on similar prompts. Concretely, we showed that Meta-Llama-3.1-8B-Instruct can be converted into a CLM on par with or exceeding the performance of modern CLMsmol-gen-bg3, mol-gen-bg4, mol-gen-gpt1, mol-opt1, iminer, rather than only demonstrating that LLMs can modify currently existing SMILES strings to have different properties.
Our study clarifies a few crucial points for CLMs derived from LLMs going forward. First, it is not necessary to pretrain a specialized model on chemistry-specific text to generate molecules from a text description; a much less resource-intensive SFT training run on prompt-following on a dataset of a few million molecules with a commodity LLM is enough to do that. Second, DPO provides another resource-efficient way of optimizing the model to produce molecules that score well on a targeted objective without needing in-context examples, instead relying on the generative nature of the model itself for good and bad examples. Using these optimization methods, we showed that data specific to a target protein can generate increasingly better inhibitor molecules as determined by docking scores. A corollary to this is the finding that SmileyLlama can combine its knowledge gained during single-objective optimization to perform well at a task specifying multiple objectives, elicited by combining the prompts, which is a welcome outcome. SmileyLlama was able to generate molecules which satisfy Lipinski’s rule of 5 and a high docking score on SARS2-Mpro even though SmileyLlama was not trained on the generation of molecules with both prompts, but instead is found to generate molecules with both attributes at inference.
It should be mentioned that both high docking and drug-likeness scores are not enough for a molecule to be a good drug candidate; good candidates must also inhibit so-called "off-target effects", must be robust to mutation of the virus (generally through binding to essential residues), among other downstream requirements. While SmileyLlama is not being explicitly optimized for generating molecules with these qualities in this work, the DPO framework laid out here is extensible to optimizing molecules for these characteristics: for instance, off-target effects can be avoided by penalizing docking to proteins similar to the target. Furthermore, there are still limitations to and tradeoffs within the SmileyLlama framework and our investigation. While DPO improves adherence to the prompt, it does so at the cost of narrowing the distribution of properties or diversity, which may not be desirable in all application areas outside of discovery. Furthermore, SmileyLlama still struggles in data-poor regimes, for instance in the task of generating macrocycles.
The prompting and optimization framework for modifying LLMs to explore chemical space broadly or to narrow the search to specific regions shown here could also be leveraged for molecular design outside of drug discovery, such as the use of SMILES for elaborating on transition metal complexesRasmussen2025. One could also imagine that casting a chemical problem as a linguistic construct could also allow for new applications, such as our recent work on chemical synthesisSun2025. As with many of the fields touched by LLMs this decade, the newly opened frontier of possibility in chemistry is as vast as it is exciting.
Acknowledgments
This work was supported in part by the National Institute of Allergy and Infectious Disease grant U19-AI171954 for the drug molecule application. We thank the CPIMS program, Office of Science, Office of Basic Energy Sciences, Chemical Sciences Division of the U.S. Department of Energy under Contract DE-AC02-05CH11231 for support of the machine learning. We thank Riza Özçelik for kindly providing the retraining code for different CLMs for benchmarking and Nicole Kennedy for suggesting properties of molecules useful to medicinal chemists.
Data Availability
The data included here contains all SMILES strings we used from ChEMBL (chembl.txt), the dataset containing the BRICS substructures of these ChEMBL molecules, the dataset containing prompts and responses used for SFT to create SmileyLlama from Llama, and the dataset used to create SmileyLlama-DPO-opt from SmileyLlama. Source data for all Figures is available with this manuscript. Data used for this study can be found at https://anonymous.4open.science/r/SmileyLlama-gh-D038.
Code Availability
Code used for this study can be found at https://anonymous.4open.science/r/SmileyLlama-gh-D038/. SmileyLlama and its derivatives optimized with DPO can be found at https://huggingface.co/collections/ THGLab/smileyllama-6880b14e4e4d708001564062
Author Contributions Statement
J.M.C.,K.S., A.G., and T.H.G. defined goals and designed the project. J.M.C., K.S., A.G., and D.B. carried out the optimizations. J.M.C., K.S., A.G., and T.H.G. wrote the paper. All authors discussed the results and made comments and edits to the manuscript.
Competing Interests Statement
The authors declare no competing interests.